OpenStudy (anonymous):

Please help! If the center is at (-3,-1), the horizontal semi-major axis 7 units and the semi-minor axis 5 units long what is the equation of the ellipse?

OpenStudy (anonymous):

\[\frac{ (x+3)^{2} }{ 5 }+\frac{( y+1)^{2} }{ 7 }=1\] Is this right?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

Hey, AGAIN, square the denominator, PLEASE One more thing: you have "horizontal major" that means your ellipse is the form 

OpenStudy (loser66):

7>5 so a will be 7, and b =5

OpenStudy (loser66):

Your formula should be \[\dfrac{x+3)^2}{7^2}+\dfrac{(y+1)^2}{5^2}=1\] I repeat it again. SQUARE the denominator.

OpenStudy (loser66):

Got me?

OpenStudy (anonymous):

Thank you i thought this problem was solved a bit differently, guess not, thanks!

OpenStudy (loser66):

I know, It is so complicated and we have to remember all. That's why I am here to re-study everything by helping others. I am not good. Every time I help you, I re-read my note. After 100 times, I remember the cases. hahahaha.....